# Is the type-D NUT C-metric really “missing” from the most general Plebański-Demiański solution?

### v1

@@ Abstract
It remains a long-standing problem, unsettled for almost two decades in General Relativity community, ever since Griffiths and Podolsky demonstrated in their previous paper [J.B. Griffiths and J. Podolsky, Classical Quantum Gravity 22, 3467 (2005)] that the type-D NUT C-metric seems to be absent from the most general family of the type-D Plebański-Demiański (P-D) solution. However, Astorino, in his recent article [M. Astorino, Phys. Rev. D 109, 084038 (2024)] presented a different form of rotating and accelerating black holes and showed that all known four-dimensional type-D accelerating black holes (without the NUT charge) can be recovered via various different limits in a definitive fashion. In particular, he provided, for the first time, the correct expressions for the type-D static accelerating black holes with a nonzero NUT charge, which was previously impossible using the traditional parametrization of the familiar P-D solution. Nevertheless, it still remains elusive that how these two different forms of the four-dimensional rotating and accelerating solutions are related. In this paper, we aim to fill this gap by finding the obvious coordinate transformations and parameter identifications between the vacuum metrics after two different parameterizations of the generated solution via the inverse scattering method from the seed metric – the Rindler vacuum background. We then resolve this “missing” puzzle by providing another Möbius transformation and linear combinations of the Killing coordinates, which clearly cast the type-D NUT C-metric into the familiar form of the P-D solution. Additionally, we propose an alternative new routine for the normalization of the obtained metric derived via the inverse scattering method from the vacuum seed solution, which could be potentially useful for the construction of higher-dimensional solutions using the trivial vacuum background as the seed metric.
@@ I Introduction
The most general type-D solution in General Relativity has been well-known for more than half a century JMP10-1195 ; AP90-196 ; AP98-98 , and is often referred to as the Plebański-Demiański (P-D) metric. With the advent of the conveniently factorized form CQG22-109 of two quartic structure functions, this solution has subsequently been extensively studied during the past twenty years, see the book GP2009 for a comprehensive review. However, as demonstrated in Ref. CQG22-3467 , the type-D NUT C-metric is apparently accelerating and can be transformed into the usual Taub-NUT solution. Since then, “such a solution has to date neither been identified nor proved not to exist.” This raised a “missing” puzzle for almost two decades, as the type-D NUT C-metric appears to be absent from the most general family of the type-D P-D solution.
@@
Recently, Astorino PRD109-084038 presented a different form of the rotating and accelerating black holes that belong to the type-D family (although no obvious proof of this issue is provided in that paper) and showed that all known four-dimensional type-D accelerating black holes without the NUT charge can be obtained via various different limits in a definitive fashion. Moreover, he presented, for the first time, the correct expressions for the type-D NUT C-metric, which are not available in the traditional parametrization for the familiar P-D solution. Lately, this solution has also been generalized in 2404.06551 to include the electro-magnetic charges and a nonzero cosmological constant.
@@
As far as the NUT C-metric is concerned, apart from the above type-D one, there is also another kind of the accelerating Taub-NUT solution PRD74-084031 , which belongs to a more general type-I PRD102-084024 that has attracted some recent interest PRD108-124025 ; PRD108-024059 ; 2309.13656 . Although the initial form of this solution is complicated and has been obtained via a hybrid solution-generating method from a specific seed, in fact it can be simply generated PRD108-124025 ; PRD108-024059 via the Ehler transformation from the C-metric. The Ehler transformation not only endows the black hole horizon with the NUT charge, but also allocates a port of the NUT charge to the accelerating horizon. From the perspective of the double Kerr-Schild formalism, the type-D NUT C-metric can be viewed as a perturbation on the pure Rindler background. It seems that one could regard the NUT-charged Rindler metric as the background spacetime of the type-I NUT C-metric. However, it is unclear whether the type-I C-metric admits a similar double Kerr-Schild formulation. Based on the same reasoning, when the Harrison transformation is applied to the seed C-metric, one gets a type-I electrically charged C-metric, whose background is the electrified Rindler vacuum. However, a well-known type-D charged static accelerating solution (Reissner-Nordström C-metric, or RN C-metric) is already included in the most general P-D family. Like the Ehler transformation that allocates a port of the NUT charge to the Rindler background, the Harrison transformation also assigns a port of the electric charge to the Rindler background. Thus, in the type-I RN C-metric, the accelerating horizon acquires the electric charge; in other words, its background is a charged Rindler metric. In this sense, the original Kinnersley-Walker’s interpretation PRD2-1359 of the uniformly accelerating charged mass as two “oppositely” accelerated black holes apart from each other seems not to be very accurate, as it can not distinguish the type-D accelerating black holes from the type-I ones. Rather, one could say that the type-D accelerating solution is a black hole siting on the pure Rindler background, while the type-I accelerating metric is that on the more general Rindler background with dynamical parameters (the NUT charge, electro-magnetic charge) plus other kinetic parameters (such as the rotation parameter and possibly the cosmological constant). It is noted that while the type-D solution admits an extension to include a nonzero cosmological constant, the type-I metric seems unable to permit such a generalization. By the way, we mention that this is also another kind of the NUT C-metric presented in Ref. PRD91-124005 , which is dubbed the [4-1]-soliton solution PRD93-044021 . This solution still remains unexplored to date, but we can anticipate it to be of type-I in the Petrov classification.
@@
What remains opaque in the Astorino’s work is how these two different forms of the four-dimensional rotating and accelerating solutions are related, and why the type-D NUT C-metric is “missing” from the P-D metric. The main purpose of this paper is to resolve this “missing” puzzle in two different ways. First, after generating the rotating and accelerating Kerr-NUT metric via the inverse scattering method (ISM) JETP48-985 ; JETP50-1 ; BV2005 from the seed Rindler vacuum background, we obtain two distinct forms of the same solution by using two different parameterizations. This clearly shows that both metrics are isometric to each other. Then, we find the obvious Möbius transformations and parameter identifications between them. In the meanwhile, we will illustrate what limit should be taken in the general P-D solution to obtain the type-D NUT C-metric. In particular, we shall present another Möbius transformation and linear combinations of the Killing coordinates that clearly cast the type-D NUT C-metric into the familiar form of the P-D solution, with the quartic structure function being factorized into two quadratic functions in a manner different from that of Griffiths-Podolsky’s.
@@
The main parts of our paper is organized as follows. In Sec. II, we present the ISM construction of the type-D accelerating Kerr-NUT metric as a 2-soliton transformation from the Rindler vacuum background, and two distinct parameterizations to arrive at two different rational expressions of the same solution. As preparation, we take the C-metric as a simple example to compare the double Kerr-Schild formalism with the ISM perspective, and attempt to combine these two seemly unrelated schemes. This idea has never been advocated in the past years. Then after a minor retrospect of the ISM algorithm, we devise a new strategy for the normalization of the obtained metric derived via the ISM construction from the vacuum seed solution, which could facilitate the construction of higher dimensional solutions by using the trivial vacuum background as the seed metric. Our normalization scheme is corroborated in the Appendix Appendix: ISM construction of 5-dimensional static solutions with our new normalization scheme by deriving the 5-dimensional Schwarzschild-Tangherlini solution NC27-636 from the flat vacuum background in two different ways. The remaining two sections III and IV are entirely new and constitute the main contribution of this work, where Möbious transformations of the radial and angular coordinates, and linear combinations of the Killing coordinates as well as various parameter identifications between these two metrics are plainly displayed. Our paper is ended up with a brief summary of this work, followed by the appendix with two simple examples to illuminate our new normalization strategy.
II Accelerating Kerr-NUT solution: ISM construction and two different parameterizations
@@ II.1 The C-metric and its seed
We begin with two different descriptions of the C-metric, whose line element in spherical-like coordinates ($t,r,x=\cos\theta,\bar{\phi}$)
@@
in which two structure functions being
@@
where the 3-soliton $\mu_{k}$’s are algebraically expressed as
@@
which will be used later as our seed metric. The Rindler metric can be viewed as 1-soliton transformation from the flat Minkowski background with the metric components: $g_{\mu\nu}=(-1,\rho^{2},1,1)$. However, it is noted that the Rindler vacuum is trivial and is entirely equivalent to flat Minkowski metric since they are related via some simple coordinate transformations.
So in the soliton formalism, the C-metric can be interpreted either as a 3-soliton transformation to the flat Minkowski background or a 2-soliton transformation to the Rindler background. In this article, we will take the latter viewpoint and derive the accelerating Kerr-NUT solution via the ISM procedure by using the Rindler background (6) as the seed metric. Note that usually the seed solution need not to be exactly the background metric and frequently can be taken as more complicated even singular one for the technical aim. However, a simple background metric is surely more preferable in practice as a seed solution.
@@
$+\frac{2m\big{(}1-\alpha^{2}r^{2}\big{)}}{r(1+\alpha\,rx)^{2}}\Big{(}d\bar{t}+\frac{dr}{1-\alpha^{2}r^{2}}\Big{)}^{2}+\frac{2m\alpha\,x\big{(}1-x^{2}\big{)}r^{2}}{(1+\alpha\,rx)^{2}}\Big{(}d\bar{\phi}+\frac{I\,dx}{1-x^{2}}\Big{)}^{2}\,.$
@@
However, it seems that the following double quasi-Kerr-Schild prescription
@@
is in more accordance with the above 2-soliton viewpoint by virtue of the 2-dimensional metrics are conformal to each other:
@@ II.2 ISM: A minor preliminary
In a $D$-dimensional space-time with $(D-2)$ Killing vectors, if its metric, whose components depend on the Weyl coordinates $\rho$ and $z$ only, is written as PRD70-124002
@@
and supplemented with the canonical condition:
@@
in which $\mathbf{n}^{{\dagger}(k)}$ denotes the transpose of the BZ vector $\mathbf{n}^{(k)}$ defined below, $\mu_{k}$ represents a soliton with its corresponding anti-soliton being $\nu_{k}=-\rho^{2}/\mu_{k}$, and the constant $z_{k}$ is the turning point (or rod endpoint) on the $z$ axes.
The expression for the generated matrix of the general $n$-soliton solutions is:
$\widetilde{\mathbf{g}}=\mathbf{g}_{0}-\sum_{k,l=1}^{n}\frac{\mathbf{l}^{{\dagger}(k)}(\mathbf{\Gamma}^{-1})_{kl}\mathbf{l}^{(l)}}{\mu_{k}\mu_{l}}\,.$ (16)
@@
In the case of a diagonal seed, the generating matrix $\mathbf{\Psi}_{0}(\mu_{i},\rho,z)$ can be easily obtained from the seed metric $\mathbf{g}_{0}(\rho,z)$ by simply replacing every soliton or anti-soliton via $\mu_{k}\to\mu_{k}-\lambda$ and $\rho^{2}\to\rho^{2}-2\lambda\,z-\lambda^{2}$, or $\rho^{2}/\mu_{k}\to\rho^{2}/\mu_{k}+\lambda$.
@@
As far as the normalization scheme is concerned, to the best of our knowledge, there exist three different kinds of strategies up to date. (I) Uniform normalization BV2005 : Like the four-dimensional case, all the metric components of the Killing parts in the higher-dimensional case are multiplied by the same normalization factor. The main defect of this method lies in that it yields some fractional singularities on the symmetry axes, so that the final solution has little physics meaning. (II) Partial block uniform normalization PRD73-064009 ; PRD73-124034 ; PRD74-104004 ; PTP116-319 , which stems from early work PRD34-2990 ; PLA136-269 : This method only applies the above uniform normalization to a part of the block diagonal matrix of the generated metric, so it is only effective in the case when not all rotation parameters are turned on, for instance, it is limited to the singly rotating case in five dimensions, and so on. (III) Pomeransky’s remove–re-add tricky PRD73-044004 : If the seed metric already satisfies the canonical condition, then after removing and readding the same number of solitions or anti-solitons, the generated metric automatically subjects to the canonical condition also. At the same time, the conformal factor can be simply computed as $f=f_{0}\det\mathbf{\Gamma}/{\det\mathbf{\Gamma}_{0}}$, where $\mathbf{\Gamma}_{0}$ can be obtained from $\mathbf{\Gamma}$ by setting the nontrivial BZ constants to zero. This is a wise strategy that is widely adopted nowadays, for example, the doubly rotating Myers-Perry solution AP172-304 is re-generated via this method from the five-dimensional Schwarzschild-Tangherlini solution NC27-636 . However, there is a fly in the ointment in that it can not work effective in the case when the seed solution is a vacuum background metric where there is no soliton that can be removed. In particular, it is claimed PRD73-044004 that the five-dimensional Schwarzschild-Tangherlini solution NC27-636 cannot be obtained as a 2-soliton solution on the flat Minkowski background.
Below we propose the fourth scheme for the normalization of the obtained metric derived from the vacuum seed solution, which would be potentially useful for the construction of higher dimensional solutions by using the trivial vacuum background as the seed metric. In the appendix, we will present two simple examples to illuminate this method. First, let’s consider a simple ansatz:
@@
by simply employing a diagonal matrix $\mathbf{W}$ with which each of its entry satisfies the Laplace equation:
@@
If the generated metric is still diagonal just as its seed metric, then the above means amounts to $\mathbf{g}=\mathbf{W}^{2}\widetilde{\mathbf{g}}$, and it is easy to compute the conformal factor. On the other hand, if all the diagonal entries are identical $w_{i}=w$, then the above scheme is equivalent to the original uniform normalization strategy. Supposed that only some of the diagonal entries are set to be identical, then this amounts to the partial block uniform normalization. Finally, if all $w_{i}$’s are appropriately chosen such as ($-\rho^{2}/\mu_{i}^{2}$) or ($-\mu_{i}^{2}/\rho^{2}$), this scheme is actually equivalent to removing solition or anti-soliton at least in the case of the generated metric being static (diagonal).
Second, let’s handle the more general case. Since, the generated metric in general still needs to make further linear combinations of the Killing coordinates, that is, $\hat{\mathbf{g}}=\mathbf{A}^{{\dagger}}\mathbf{g}\mathbf{A}$, where the rotation matrix $\mathbf{A}$ is unimodular and its every entry is a constant. Then, we can have $\hat{\mathbf{g}}=\hat{\mathbf{W}}^{{\dagger}}\widetilde{\mathbf{g}}\hat{\mathbf{W}}$, where $\hat{\mathbf{W}}=\mathbf{W}\mathbf{A}$, $\hat{\mathbf{W}}^{{\dagger}}$ is the matrix transpose of $\hat{\mathbf{W}}$ and each of its entry obeys the above Laplace equation (19). Clearly it is a trouble thing to choose a suitable matrix $\hat{\mathbf{W}}$ so that the final metric is the expected one. However, one need not to worry about this, since the Pomeransky’s remove–re-add tricky is already very effective to generate the rotating solution, so our method just plays a nice complementary where it fails to work.

### v3

@@ Abstract
It remains a long-standing problem, unsettled for almost two decades in the general relativity community, ever since Griffiths and Podolský demonstrated in their previous paper [J.B. Griffiths and J. Podolský, Classical Quantum Gravity 22, 3467 (2005)] that the type-D NUT C-metric seems to be absent from the most general family of the type-D Plebański-Demiański (P-D) solution. However, Astorino, in his recent article [M. Astorino, Phys. Rev. D 109, 084038 (2024)] presented a different form of rotating and accelerating black holes and showed that all known four-dimensional type-D accelerating black holes (without the NUT charge) can be recovered via various different limits in a definitive fashion. In particular, he provided, for the first time, the correct expressions for the type-D static accelerating black holes with a nonzero NUT charge, which was previously impossible using the traditional parametrization of the familiar P-D solution. Nevertheless, it still remains elusive how these two different forms of the four-dimensional rotating and accelerating solutions are related. In this paper, we aim to fill this gap by finding the obvious coordinate transformations and parameter identifications between the vacuum metrics after two different parametrizations of the generated solution via the inverse scattering method from the seed metric – the Rindler vacuum background. We then resolve this “missing” puzzle by providing another Möbius transformation and linear combinations of the Killing coordinates, which clearly cast the type-D NUT C-metric into the familiar form of the P-D solution. Additionally, we propose an alternative new routine for the normalization of the obtained metric derived via the inverse scattering method from the vacuum seed solution, which could be potentially useful for the construction of higher-dimensional solutions using the trivial vacuum background as the seed metric.
@@ I Introduction
The most general type-D solution in general relativity has been well known for more than half a century JMP10-1195 ; AP90-196 ; AP98-98 , and is often referred to as the Plebański-Demiański (P-D) metric. With the advent of the conveniently factorized form CQG22-109 of two quartic structure functions, this solution has subsequently been extensively studied during the past twenty years, see the book GP2009 for a comprehensive review. However, as demonstrated in Ref. CQG22-3467 , the type-D NUT C-metric is apparently accelerating and can be transformed into the usual Taub-NUT solution. Since then, “such a solution has to date neither been identified nor proved not to exist.” This raised a “missing” puzzle for almost two decades, as the type-D NUT C-metric appears to be absent from the most general family of the type-D P-D solution.
@@
Recently, Astorino PRD109-084038 presented a different form of the rotating and accelerating black holes that belong to the type-D family (although no obvious proof of this issue is provided in that paper) and showed that all known four-dimensional type-D accelerating black holes without the NUT charge can be obtained via various different limits in a definitive fashion. Moreover, he presented, for the first time, the correct expressions for the type-D NUT C-metric, which are not available in the traditional parametrization for the familiar P-D solution. Lately, this solution has also been generalized in 2404.06551 to include the electromagnetic charges and a nonzero cosmological constant.
@@
As far as the NUT C-metric is concerned, apart from the above type-D one, there is also another kind of accelerating Taub-NUT solution PRD74-084031 , which belongs to a more general type I PRD102-084024 that has attracted some recent interest GBoldi ; PRD108-124025 ; PRD108-024059 ; JHEP0823085 ; GRG56-111 . Although the initial form of this solution is complicated and has been obtained via a hybrid solution-generating method from a specific seed, in fact it can be simply generated GBoldi ; PRD108-124025 ; PRD108-024059 via the Ehlers transformation from the C-metric. The Ehlers transformation not only endows the black hole horizon with the NUT charge, but also allocates a port of the NUT charge to the accelerating horizon. From the perspective of the double Kerr-Schild formalism, the type-D NUT C-metric can be viewed as a perturbation on the pure Rindler background. It seems that one could regard the NUT-charged Rindler metric as the background spacetime of the type-I NUT C-metric. However, it is unclear whether the type-I C-metric admits a similar double Kerr-Schild formulation. Based on the same reasoning, when the Harrison transformation is applied to the seed C-metric, one gets a type-I electrically charged C-metric, whose background is the electrified Rindler vacuum. However, a well-known type-D charged static accelerating solution (Reissner-Nordström C-metric, or RN C-metric) is already included in the most general P-D family. Like the Ehlers transformation that allocates a port of the NUT charge to the Rindler background, the Harrison transformation also assigns a port of the electric charge to the Rindler background. Thus, in the type-I RN C-metric, the accelerating horizon acquires the electric charge; in other words, its background is a charged Rindler metric. In this sense, the original Kinnersley-Walker’s interpretation PRD2-1359 of the uniformly accelerating charged mass as two “oppositely” accelerated black holes apart from each other seems not to be very accurate, as it cannot distinguish the type-D accelerating black holes from the type-I ones. Rather, one could say that the type-D accelerating solution is a black hole sitting on the pure Rindler background, while the type-I accelerating metric is that on the more general Rindler background with dynamical parameters (the NUT charge, electromagnetic charge) plus other kinetic parameters (such as the rotation parameter and possibly the cosmological constant). It is noted that while the type-D solution admits an extension to include a nonzero cosmological constant, the type-I metric seems unable to permit such a generalization. By the way, we mention that there is also another kind of NUT C-metric presented in Ref. PRD91-124005 , which is dubbed the [4-1]-soliton solution PRD93-044021 . This solution still remains unexplored to date, but we can anticipate it to be of type I in the Petrov classification.
@@
What remains opaque in the Astorino’s work is how these two different forms of the four-dimensional rotating and accelerating solutions are related, and why the type-D NUT C-metric is “missing” from the P-D metric. The main purpose of this paper is to resolve this “missing” puzzle in two different ways. First, after generating the rotating and accelerating Kerr-NUT metric via the inverse scattering method (ISM) JETP48-985 ; JETP50-1 ; BV2005 from the seed Rindler vacuum background, we obtain two distinct forms of the same solution by using two different parametrizations. This clearly shows that both metrics are isometric to each other. Then, we find the obvious Möbius transformation and parameter identifications between them. In the meanwhile, we will illustrate what limit should be taken in the general P-D solution to obtain the type-D NUT C-metric. In particular, we shall present another Möbius transformation and linear combinations of the Killing coordinates that clearly cast the type-D NUT C-metric into the familiar form of the P-D solution, with the quartic structure function being factorized into two quadratic functions in a manner different from that of Griffiths-Podolský.
@@
The main parts of our paper are organized as follows. In Sec. II, we present the ISM construction of the type-D accelerating Kerr-NUT metric as a 2-soliton transformation from the Rindler vacuum background, and two distinct parametrizations to arrive at two different rational expressions of the same solution. As preparation, we take the C-metric as a simple example to compare the double Kerr-Schild formalism with the ISM perspective, and make an attempt to combine these two seemly unrelated schemes. This idea has never been advocated in the past years. Then after a minor retrospect of the ISM algorithm, we devise a new strategy for the normalization of the obtained metric derived via the ISM construction from the vacuum seed solution, which could facilitate the construction of higher dimensional solutions by using the trivial vacuum background as the seed metric. Our normalization scheme is corroborated in the Appendix Appendix: ISM construction of five-dimensional static solutions with our new normalization scheme by deriving the five-dimensional Schwarzschild-Tangherlini solution NC27-636 from the flat vacuum background in two different ways. The remaining two sections, III and IV, are entirely new and constitute the main contribution of this work, where the Möbious transformation of the radial and angular coordinates, and linear combinations of the Killing coordinates as well as various parameter identifications between these two metrics are plainly displayed. Our paper ends with a brief summary of this work, followed by the Aappendix with two simple examples to illuminate our new normalization strategy.
II Accelerating Kerr-NUT solution: ISM construction and two different parametrizations
@@ II.1 The C-metric and its seed
We begin with two different descriptions of the C-metric, whose line element in the spherical-like coordinates ($t,r,x=\cos\theta,\bar{\phi}$),
@@
with two structure functions being
@@
where the 3-soliton $\mu_{k}$s are algebraically expressed as
@@
which will be used later as our seed metric. The Rindler metric can be viewed as 1-soliton transformation from the flat Minkowski background with the metric components: $g_{\mu\nu}=(-1,\rho^{2},1,1)$. However, it is noted that the Rindler vacuum is trivial and is entirely equivalent to the flat Minkowski metric since they are related via some simple coordinate transformations.
So in the soliton formalism, the C-metric can be interpreted either as a 3-soliton transformation to the flat Minkowski background or a 2-soliton transformation to the Rindler background. In this article, we will take the latter viewpoint and derive the accelerating Kerr-NUT solution via the ISM procedure by using the Rindler background (6) as the seed metric. Note that usually the seed solution need not be exactly the background metric and frequently can be taken as a more complicated even singular one for the technical aim. However, a simple background metric is surely more preferable in practice as a seed solution.
@@
$+\frac{2m\big{(}1-\alpha^{2}r^{2}\big{)}}{r(1+\alpha\,rx)^{2}}\Big{(}d\bar{t}+\frac{dr}{1-\alpha^{2}r^{2}}\Big{)}^{2}+\frac{2m\alpha\,x\big{(}1-x^{2}\big{)}r^{2}}{(1+\alpha\,rx)^{2}}\Big{(}d\bar{\phi}+\frac{I\,dx}{1-x^{2}}\Big{)}^{2}\,,$
@@
However, it seems that the following double quasi-Kerr-Schild prescription,
@@
is in more accordance with the above 2-soliton viewpoint by virtue of the two-dimensional metrics are conformal to each other:
@@ II.2 ISM: A minor preliminary
In a $D$-dimensional spacetime with $(D-2)$ Killing vectors, if its metric, whose components depend on the Weyl coordinates $\rho$ and $z$ only, is written as PRD70-124002
@@
and supplemented with the canonical condition,
@@
in which $\mathbf{n}^{{\dagger}(k)}$ denotes the transpose of the BZ vector $\mathbf{n}^{(k)}$ defined below, $\mu_{k}$ represents a soliton with its corresponding antisoliton being $\nu_{k}=-\rho^{2}/\mu_{k}$, and the constant $z_{k}$ is the turning point (or rod end point) on the $z$ axes.
The expression for the generated matrix of the general $n$-soliton solutions is
$\widetilde{\mathbf{g}}=\mathbf{g}_{0}-\sum_{k,l=1}^{n}\frac{\mathbf{l}^{{\dagger}(k)}(\mathbf{\Gamma}^{-1})_{kl}\mathbf{l}^{(l)}}{\mu_{k}\mu_{l}}\,,$ (16)
@@
In the case of a diagonal seed, the generating matrix $\mathbf{\Psi}_{0}(\mu_{i},\rho,z)$ can be easily obtained from the seed metric $\mathbf{g}_{0}(\rho,z)$ by simply replacing every soliton or antisoliton via $\mu_{k}\to\mu_{k}-\lambda$ and $\rho^{2}\to\rho^{2}-2\lambda\,z-\lambda^{2}$, or $\rho^{2}/\mu_{k}\to\rho^{2}/\mu_{k}+\lambda$.
@@
As far as the normalization scheme is concerned, to the best of our knowledge, there exist three different kinds of strategies to date. (I) Uniform normalization BV2005 : Like the four-dimensional case, all the metric components of the Killing parts in the higher-dimensional case are multiplied by the same normalization factor. The main defect of this method lies in that it yields some fractional singularities on the symmetry axes, so that the final solution has little physics meaning. (II) Partial block uniform normalization PRD73-064009 ; PRD73-124034 ; PRD74-104004 ; PTP116-319 , which stems from early work PRD34-2990 ; PLA136-269 : This method only applies the above uniform normalization to a part of the block diagonal matrix of the generated metric, so it is only effective in the case when not all rotation parameters are turned on, for instance, it is limited to the singly rotating case in five dimensions, and so on. (III) Pomeransky’s remove–re-add tricky PRD73-044004 : If the seed metric already satisfies the canonical condition, then after removing and re-adding the same number of solitions or antisolitons, the generated metric automatically subjects to the canonical condition also. At the same time, the conformal factor can be simply computed as $f=f_{0}\det\mathbf{\Gamma}/{\det\mathbf{\Gamma}_{0}}$, where $\mathbf{\Gamma}_{0}$ can be obtained from $\mathbf{\Gamma}$ by setting the nontrivial BZ constants to zero. This is a wise strategy that is widely adopted nowadays, for example, the doubly rotating Myers-Perry solution AP172-304 is regenerated via this method from the five-dimensional Schwarzschild-Tangherlini solution NC27-636 . However, there is a fly in the ointment in that it cannot work effectively in the case when the seed solution is a vacuum background metric where there is no soliton that can be removed. In particular, it is claimed PRD73-044004 that the five-dimensional Schwarzschild-Tangherlini solution NC27-636 cannot be obtained as a 2-soliton solution on the flat Minkowski background.
Below we propose the fourth scheme for the normalization of the obtained metric derived from the vacuum seed solution, which would be potentially useful for the construction of higher dimensional solutions by using the trivial vacuum background as the seed metric. In the Appendix, we will present two simple examples to illuminate this method. First, let us consider a simple ansatz:
@@
by simply employing a diagonal matrix $\mathbf{W}$ with which each of its entries satisfies the Laplace equation:
@@
If the generated metric is still diagonal just as its seed metric, then the above means amounts to $\mathbf{g}=\mathbf{W}^{2}\widetilde{\mathbf{g}}$, and it is easy to compute the conformal factor. On the other hand, if all the diagonal entries are identical $w_{i}=w$, then the above scheme is equivalent to the original uniform normalization strategy. Suppose that only some of the diagonal entries are set to be identical, then this amounts to the partial block uniform normalization. Finally, if all $w_{i}$s are appropriately chosen such as ($-\rho^{2}/\mu_{i}^{2}$) or ($-\mu_{i}^{2}/\rho^{2}$), this scheme is actually equivalent to removing solition or antisoliton at least in the case of the generated metric being static (diagonal).
Second, let us handle the more general rotating case. Since the generated metric in general still needs to make further linear combinations of the Killing coordinates, that is, $\hat{\mathbf{g}}=\mathbf{A}^{{\dagger}}\mathbf{g}\mathbf{A}$, where the rotation matrix $\mathbf{A}$ is unimodular and its every entry is a constant, then we can have $\hat{\mathbf{g}}=\hat{\mathbf{W}}^{{\dagger}}\widetilde{\mathbf{g}}\hat{\mathbf{W}}$, where $\hat{\mathbf{W}}=\mathbf{W}\mathbf{A}$, $\hat{\mathbf{W}}^{{\dagger}}$ is the matrix transpose of $\hat{\mathbf{W}}$ and each of its entry obeys the above Laplace equation (19). Clearly it is troublesome to choose a suitable matrix $\hat{\mathbf{W}}$ so that the final metric is the expected one. However, one need not worry about this, since the Pomeransky’s remove–re-add tricky is already very effective to generate the rotating solution, so our method just plays a nice complement to the case where the Pomeransky’s tricky fails to work.